THE SPECTROSCOPY STUDY OF UF AND UF^{+}

\underline{J. H. BARTLETT}, I. O. ANTONOV, M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Uranium metal has been laser ablated in a pulsed-jet expansion of 0.1% SF_{6} in He to generate the fluoride. Spectroscopic data are obtained using the laser-induced fluorescence and resonance-enhanced multiphoton ionization techniques for gas-phase UF for the first time. The pulsed-field ionization zero electron kinetic energy (PFI-ZEKE) technique is used to study the low-lying states of UF^{+}. The ionization potential from PFI-ZEKE for UF is found to be D_{e} = 51137 cm^{−1}. Ω states up to \frac{13}{2} and vibrational states v^{+} = 0-4 are observed in the REMPI and PFI-ZEKE spectra. Assignments of the electronically excited states are based on measured Ω values and high-level ab initio calculations.